# Multi-scale optimal control for Einstein Telescope active seismic isolation

### v1

@@ 3.3 Cost function
Where $\mathcal{K}$ is the set of all possible filters which satisfy extra constraints which could be enforced for practical reasons. For example, In this work, we need to have a finite upper Unity Gain Frequency (UGF) for both loops. This translates to the following constraint for $\theta$ loop:
@@
The $\gamma_{1}$ is the maximum deviation from the $\xi_{\theta}$ — which measures how much we are willing to deviate from the acausal optimum. The smaller the $\gamma_{1}$ the better the performance on $\theta$ but this would restrict the performance on the x loop. Since we don’t have much constraint on the $\theta$ loop except the finite UGF we can set this value in a relaxed way. Furthermore, instead of enforcing an $H_{\infty}$  norm on the $\theta$  loop, it is possible to constrain the unity gain frequency (UGF) by introducing a right-half-plane zero paired with a left-half-plane pole at similar frequency. This approach shapes the open-loop transfer function to enforce a desired roll-off behavior.
@@ 4 Results
Table 1: Optimization results for both OmniSens and T360+BRS configurations. The reported values indicate the number of states in each system (i.e., the size of the state-space $A$ matrix). $P$ denotes the generalized plant for each sensor setup. $K_{x}$ and $K_{\theta}$ represent the number of states in the feedback controllers for the $x$ and $\theta$  loops, respectively, while $H_{x}$ and $H_{\theta}$ correspond to the number of states in the blending filters (including both high-pass and low-pass components).
@@ A.1 Definitions
$p_{l}$ pendulum’s length 0.7 $m$
$p_{d}$ pendulum’s second part length $-1.8\times 10^{-4}$ $m$
@@
Table 2: Variables and parameters used in the model. (update these parameters)
@@ A.2 Equation of Motion
Figure 12: $dy2-dy1$ (vertical sensors) response to input translation $x_{t}$ and rotation $\phi_{t}$. dx (horizontal sensor) response to input translation $x_{t}$ and rotation $\phi_{t}$
@@ A.3 Thermal noise in the reference mass
The thermal noise would contribute to the rotational performance of the pendulum. It would act as a torque on the pendulum which could be modelled as a sensing noise. The power spectral density of the thermal noise is given by [15]:
@@
Where $K_{B}$ is the Boltzmann constant, $T$ is the room temperature, $\phi_{eff}$ is the effective loss angle of the pendulum and $k_{el}$ is the spring constant. This noise will excite the center of mass of pendulum and result in rotational motion which will be seen by the sensors. To model this we quadrature sum the thermal noise with the sensor noise. This will then inverted back to $x_{t}$ and $\phi_{t}$ to get the inertial equivalent noise of the reference mass which will be used in the modeling of the OmniSens configuration.
@@ Appendix B ISI parameters
[…]
@@
The rotational stiffness terms are solved in Eqs. (36) - (38).
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix C Plant transfer function matrix
The generalized plant maps $\vec{u}$ to $\vec{y}$ where these vectors are defined as:
@@
The transfer function from input to output of generalized plant is given by:
@@
Where $W$ and $P$ are the weighting function and plant transfer function respectively. The $\odot$ is the element wise multiplication.

### v2

@@ 3.3 Cost function
Where $\mathcal{K}$ is the set of all possible filters which satisfy extra constraints which could be enforced for practical reasons. For example, In this work, we need to have a finite upper Unity Gain Frequency (UGF) for both loops. Since the sensor noise exceeds the ground-induced platform motion above a certain frequency, the acausal optimum necessarily coincides with the ground motion. Consequently, to ensure that the closed-loop residual motion on $\theta$ follows the acausal optimum, the controller must remain disengaged in this frequency range; otherwise, the sensing noise will be injected into the loop, increasing the residual motion beyond the acausal optimum. To enforce a residual motion that closely follows the acausal optimum, we impose the following $H_{\infty}$ constraint on the $\theta$ loop:
@@
The $\gamma_{1}$ is the maximum deviation from the $\xi_{\theta}$ — which measures how much we are willing to deviate from the acausal optimum. The smaller the $\gamma_{1}$ the better the performance on $\theta$ but this would restrict the performance on the x loop. Since we don’t have much constraint on the $\theta$ loop except the finite UGF we can set this value in a relaxed way. Furthermore, instead of enforcing an $H_{\infty}$ norm on the $\theta$ loop, it is possible to constrain the unity gain frequency (UGF) by introducing a right-half-plane zero paired with a left-half-plane pole at a similar frequency. These add extra phase loss at high frequencies, enforcing lower unity gain. This approach shapes the open-loop transfer function to enforce a desired roll-off behavior.
@@ 4 Results
Table 2: Optimization results for both OmniSens and T360+BRS configurations. The reported values indicate the number of states in each system (i.e., the size of the state-space AA matrix). $P$ denotes the generalized plant for each sensor setup. $K_{x}$ and $K_{\theta}$ represent the number of states in the feedback controllers for the $x$ and $\theta$ loops, respectively, while $H_{x}$ and $H_{\theta}$ correspond to the number of states in the blending filters (including both high-pass and low-pass components).
@@ A.1 Definitions
$p_{l}$ pendulum’s length 0.7 mm
$p_{d}$ pendulum’s second part length $1.8\times 10^{-4}$ $m$
@@
Table 3: Variables and parameters used in the model.
@@ A.2 Equation of Motion
Figure 12: (a,b) dx (horizontal sensor) response to input translation $x_{t}$ and rotation $\phi_{t}$ . (c,d) dy2dy1dy2-dy1 (vertical sensors) response to input translation $x_{t}$ and rotation $\phi_{t}$.
@@ A.3 Thermal noise in the reference mass
The thermal noise would contribute to the rotational performance of the pendulum. It acts as a torque on the pendulum which could be modelled as a sensing noise. The power spectral density of the thermal noise is given by [9]:
@@
Where $K_{B}$ is the Boltzmann constant, $T$ is the room temperature, $\phi_{eff}$ is the effective loss angle of the pendulum and $k_{el}$ is the spring constant. This noise will excite the center of mass of pendulum and result in rotational motion which will be seen by the sensors. To model this we quadrature sum the thermal noise with the sensor noise. This will then be inverted back to $x_{t}$ and $\phi_{t}$ to get the inertial equivalent noise of the reference mass which will be used in the modeling of the OmniSens configuration.
@@ Appendix B ISI parameters
The Internal Seismic Isolation (ISI) for OmniSens is an adapted design of the LIGO Horizontal Access Module ISI (HAM-ISI), as found in [28]. The ISI is a single-stage actively and passively isolated platform. Passive isolation is provided through the suspension, which employs three blade springs (design adapted from [28]) and flexures wires/pendula (design adapted from [28]), manufactured from Titanium Grade-19. Active isolation is achieved through six actuators. The isolated platform is a bolted aluminum rigid hexagonal structure with an optical bench, has a moving mass of $415.35\>\text{kg}$, and a total mass of $625\>\text{kg}$ including payload.
@@
Where here $M$ is the mass matrix, $K$ stiffness matrix, and $u$ is the center of mass position. All parameters are defined in the table 4. The rotational stiffness terms are solved in Eqs. (34) - (36).
$k_{\phi}=$ $\frac{3}{2}k_{zz}r_{s}^{2}+3k_{yy}h^{2}-m_{u}gh-m_{s}gh_{s}\quad\text{rotation around $x$-axis}$ (34)
$k_{\theta}=$ $\frac{3}{2}k_{zz}r_{s}^{2}+3k_{xx}h^{2}-m_{u}gh-m_{s}gh_{s}\quad\text{rotation around $y$-axis}$ (35)
$k_{\psi}=$ $\frac{3}{\sqrt{2}}\sqrt{k_{xx}^{2}+k_{yy}^{2}}r_{s}^{2}+3k_{rz}\quad\text{rotation around $z$-axis}$ (36)
@@
Platform $m_{\text{u}}$ $415.35\>\text{kg}$ Moving mass of platform
$m_{\text{s}}$ $209.65\>\text{kg}$ Moving mass of payload
$m$ $625\>\text{kg}$ Total mass of platform
$z_{\text{CoG}}$ $694.9\>\text{mm}$ $z$-coordinate of platform center of gravity w.r.t. ground
@@
$r_{s}$ $300\>\text{mm}$ Radius of blade spring tips within platform
@@
Blade spring $k_{zz}$ $3.34\cdot 10^{4}\>\text{N/m}$ Blade spring translational stiffness along $z$-axis, obtained from FEA
Flexure $z_{\text{LZMP}}$ $612.8\>\text{mm}$ $z$-coordinate of LZMP w.r.t. ground
$k_{xx}$ $1.37\cdot 10^{4}\>\text{N/m}$ Flexure translational stiffness along $x$-axis, obtained from FEA
$k_{yy}$ $1.37\cdot 10^{4}\>\text{N/m}$ Flexure translational stiffness along $y$-axis, obtained from FEA
@@ Appendix C Plant transfer function matrix
The generalized plant maps input vector $\vec{u}$ to output $\vec{y}$ where these vectors are defined as:
@@
Where both $u$ and $y$ elements are shown in 6. The weighting functions $w_{\theta,p}$ and $w_{x,p}$ can be factored out of generalized plant for cleaner matrix representation:
@@
where $W$ and $P$ denote the weighting function and plant transfer function matrices, respectively. The $\odot$ is the element wise multiplication. The matrices $W$ and $\tilde{P}$ are given by: